OpenStudy (anonymous):

i forgot how to do this..can someone walk me through this process? QUESTION: What’s the equation of a line that passes through points (0, -1) and (2, 3)?

OpenStudy (anonymous):

y = 2x − 1 y = 2x + 1 x = 2y − 1 x = 2y + 1 i need to remember how to get it like this..

OpenStudy (anonymous):

alright just a moment lets see if i remember

OpenStudy (anonymous):

wait the exact equation i need is the 1st one you gave me? @stgreen

OpenStudy (anonymous):

i give up i dont get it.... i stop when i got to the point of (y-3)=(4/2)(x-2) @stgreen

OpenStudy (anonymous):

sooo 2(y-3)=2(x-2)?

OpenStudy (anonymous):

is it y=2x-1?

OpenStudy (anonymous):

@stgreen

OpenStudy (anonymous):

ok yeay.. but my question is.. thats the exact equation for it? so i have to remember the equation for me to able to do this kind of problem right? theres no easier way?

OpenStudy (anonymous):

do you know a site?